On January 2, 2015, Marin Corporation issued $2,150,000 of 10% bonds at 98 due December 31, 2024. Interest on the bonds is payable annually each December 31. The discount on the bonds is also being amortized on a straight-line basis over the 10 years. (Straight-line is not materially different in effect from the preferable “interest method.”)

The bonds are callable at 101 (i.e., at 101% of face amount), and on January 2, 2020, Marin called $1,290,000 face amount of the bonds and redeemed them.

Ignoring income taxes, compute the amount of loss, if any, to be recognized by Marin as a result of retiring the $1,290,000 of bonds in 2020. (Round answer to 0 decimal places, e.g. 38,548.)

Loss on redemption $   


Prepare the journal entry to record the redemption. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

January 2, 2020
Answer #1

Issue price of bonds = 2150000*.98 = 2107000

Discount on bonds payable = 2150000-2107000 = 43000

Unamortized discount = 43000/2 = 21500

Carrying value of bonds on January 1,2020 = 2107000+21500 = 2128500

Carrying value of 1290000 = 2128500/2150000*1290000 = 1277100

a) Loss on bond redemption = 1277100-(1290000*1.01) = -25800

b) Journal entry

Date account and explanation Debit Credit
Jan 2,2020 Bonds payable 1290000
Loss on redemption 25800
Discount on bonds payable 12900
Cash 1302900